OpenStudy (anonymous):

A company made a profit of $25,000 over a period of 5 years on an initial investment of $10,000. What is its annualized ROI?

OpenStudy (anonymous):

the gained money at 5 years = 25000-10000=15000 then the annualized rol =(15000)/(5)=5000

OpenStudy (anonymous):

whats the percent